# Imaging the Freeze-out Source and Extracting Strong Interaction Parameters in Relativistic Heavy Ion Collisions

### v2

@@ Abstract
Combining the femtoscopic interferometry and the optical deblurring algorithm, we implement a novel method to image the source formed in heavy ion collisions (HICs), while the interaction strength between the particle pair can be determined simultaneously. By applying this method to Au+Au collisions at $\sqrt{S_{\rm NN}}=200$ GeV, the spatial distribution of the emission source has been reconstructed for protons ($p$) and antiprotons ($\bar{p}$) from the respective $pp$ and $\bar{p}\bar{p}$ correlation functions. Within experimental uncertainties, protons and antiprotons share the same freeze-out distribution showing higher density in the center compared to the widely assumed Gaussian shape. The results evidence the matter-antimatter symmetry in coordinate space at the moment of freeze-out before the nucleons are fully randomized in the collisions.
I Introduction
@@
It is accepted that matter and antimatter are created in equal amount in the first moments after the Big Bang. Yet today, we observe a universe filled with matter far exceeding antimatter. This puzzling phenomenon is related to the profound symmetry breaking in nature and remains a challenge in nuclear and particle physics. On modern accelerators, heavy ion collisions (HICs) create conditions akin to those in the early universe. So after the success of observing the production of antimatter [1, 2, 3], people can further investigate the spatial evolution and decoupling of the matter and antimatter on the femtoscopic scale.
@@
The HBT intensity interferometry, invented by Hanbury-Brown and Twiss to detect the angular radius of Siris [4], has been generalized and applied extensively in nuclear physics [5, 6] to extract the temporal-spatial feature of the emission source formed in HICs by measuring the emitted particle pair correlation function (CF) at small relative momentum. In a basic picture, the CF is the result of the low-energy scattering between the particle pair emitted from the source, and carries the information of the interaction strength, defined by the scattering length $f_{0}$ and effective range $d_{0}$ between the pair, written as ($f_{0}$, $d_{0}$) in short.
@@
Thus, simultaneous inference of the source distribution and the interaction strength becomes feasible for either particle or antiparticle pairs. Based on the Lednický-Lyuboshitz (LL) model, STAR Collaboration has verified that protons and antiprotons exhibit nearly identical interaction properties, and the extracted source size parameters (Gaussian source) for $pp$ and $\bar{p}\bar{p}$ pairs were found to be similar [7]. Specifically, within experimental uncertainties, the parameters ($f_{0}$, $d_{0}$) are equivalent for $pp$ and $\bar{p}\bar{p}$ pairs, consistent with CPT symmetry. The Gaussian source assumption has become widely adopted in correlation function analysis, from which significant results have been obtained in the studies of particle interactions and source distributions [8, 9, 10, 11].
@@
However, by Gaussian parameterization it is assumed that all the participants are totally randomized in space prior to the freeze-out instant, and the initial state as well as the evolution details of the colliding system are lost. This strong assumption is unfavorable in the context of various recent studies. For two examples, i) high energy HIC has been applied to infer the initial shape of the colliding nuclei [12, 13]. ii) The Ohm’s law is suggestively broken because of the strong field and extremely short evolution time [14]. In both cases it is of importance to probe the spatial evolution of the colliding system. Thus, it is surely nontrivial to abandon the Gaussian assumption.
@@
Despite of the difficulty in solving inverse problem, the direct imaging of the source has been achieved without introducing the analytic parameterization [15, 16, 17]. More recently, the Bayessian-based deblurring method, known as Richardson-Lucy (RL) algorithm [18, 19] which was initially developed for optical imaging [20], has gained prominence in nuclear physics [21, 22, 23, 24, 25]. Particularly, Nzabahimana et al. [22] applied the RL algorithm in source imaging in HICs, highlighting the method’s effectiveness and robustness. In this letter, combining the RL deblurring algorithm and the LL model incorporating the pair interaction, we bring the realisation to solve the inverse problem of source imaging from the experimental CFs. The source distribution and ($f_{0}$, $d_{0}$) are determined simultaneously. The similarity of the source spatial distribution of $p$ and $\bar{p}$ formed in HICs has been demonstrated.
II Method description
@@
Defined as a function of the relative momentum $\mathbf{q}=\mathbf{p}_{1}-\mathbf{p}_{2}$ in the particle pair rest frame (PRF), the CF is given experimentally by $C(\mathbf{q})=\frac{P(\mathbf{p}_{1},\mathbf{p}_{2})}{P(\mathbf{p}_{1})P(\mathbf{p}_{2})}$, where $P(\mathbf{p}_{1})$ and $P(\mathbf{p}_{2})$ are the probabilities of detecting particles with momenta $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, respectively, and $P(\mathbf{p}_{1},\mathbf{p}_{2})$ is the joint probability of detecting both particles simultaneously. For uncorrelated emissions, $C=1$ is expected. Theoretically, the CF can be calculated using the Koonin-Pratt (KP) formula [26, 27]
@@
where $S$ is the source function (SF) related to the relative distance $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ of the particle-pair in the source rest frame, and $\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})$ is the wave function calculated from the relative momentum $\mathbf{q}$ and distance $\mathbf{r}^{\prime}$ in the PRF.
The LL model [28] provides a framework to describe the interaction strength between particle pairs through the parameters ($f_{0}$, $d_{0}$), which characterize the S-wave scattering amplitude under the effective range approximation. The model assumes that higher-order correlations involving three or more particles are negligible, the electromagnetic effects generated by the source are insignificant, and the momentum transfer from final-state interactions is small enough to apply the smoothness approximation. Within this model, the relative wave function for $pp$ and $\bar{p}\bar{p}$ pairs in the PRF is expressed as
$\psi_{\mathbf{q}}(\mathbf{r})$ $=e^{i\delta_{c}}\sqrt{A_{c}(\eta)}[e^{-i\frac{1}{2}\mathbf{q}\cdot\mathbf{r}}F(-i\eta,1,i(\frac{1}{2}\mathbf{q}\cdot\mathbf{r}+\frac{1}{2}qr))$ (2)
$+f_{c}(q)\frac{\widetilde{G}(\frac{1}{2}qr,\eta)}{r}].$
where $F(\alpha,1,z)$ is the confluent hypergeometric function, and $\widetilde{G}(\frac{1}{2}qr,\eta)$ combines Coulomb functions $G_{0}$ and $F_{0}$,
$\widetilde{G}(\frac{1}{2}qr,\eta)=\sqrt{A_{c}(\eta)}(G_{0}(\frac{1}{2}qr,\eta)+iF_{0}(\frac{1}{2}qr,\eta)),$ (3)
where $A_{c}(\eta)=\frac{2\pi\eta}{e^{2\pi\eta}-1}$ with the definition of $\eta=2/(qa_{0})$ (the Bohr radius $a_{0}=57.5\rm fm$). The S-wave scattering amplitude $f_{c}(q)$ is expanded as
$f_{c}(q)\approx[\frac{1}{f_{0}}+\frac{1}{8}d_{0}q^{2}-q\eta h(\eta^{-1})-i\frac{1}{2}qA_{c}(\eta)]^{-1},$ (4)
with $h(x=\eta^{-1})=\sum_{n=1}^{\infty}\frac{1}{n[(nx)^{2}+1]}-C+lnx$, where $C\approx 0.5772$ is the Euler constant. For identical particles, only considering S-wave, Eq. (1) can be written as following,
$C(\mathbf{q})=\int d^{3}rW_{d}(\mathbf{q},\mathbf{r}^{\prime})S(\mathbf{r}),$ (5)
where $W_{d}(\mathbf{q},\mathbf{r}^{\prime})=\frac{1}{2}[\frac{1}{4}|\psi_{\mathbf{q}}^{s}(\mathbf{r}^{\prime})+\psi_{\mathbf{q}}^{s}(-\mathbf{r}^{\prime})|^{2}+\frac{3}{4}|\psi_{\mathbf{q}}^{t}(\mathbf{r}^{\prime})-\psi_{\mathbf{q}}^{t}(-\mathbf{r}^{\prime})|^{2}]$.
@@
In optical deblurring, the measured distribution $f(\nu)$ can be related to the true photon distribution $\mathcal{F}(\mu)$ using a conditional probability $P(\nu|\mu)$,
@@
From the mathematical similarity between the KP formula (1) and the optical deblurring equation (6), it is naturally suggested that the RL algorithm can be applied iteratively to solve the SF. By evaluating the converged $\chi^{2}$ value, one can identify the optimal ($f_{0}$, $d_{0}$) for the wave function that best describes the CF, thereby achieving both SF imaging and extraction of strong interaction parameters simultaneously. For the convenience of numerical implementation, the KP formula is discretized by dividing $C$ into $N$ bins and $S$ into $M$ bins ($N\geq M$). The CF $C$ is expressed by $S$ through an $N\times M$ transformation matrix $K$ as
$C_{i}=\sum_{j=1}^{M}K_{ij}S_{j}.$ (7)
Once the values of $C$ and $K$ are known, the distribution of $S$ can be optimized through the RL algorithm via iterative processes:
$S_{i}^{(t+1)}=S_{i}^{(t)}\frac{\sum_{j}W_{j}\frac{C_{j}}{C_{j}^{(t)}}K_{ji}}{\sum_{j’}W_{j’}K_{j’i}},$ (8)
where $W$ represents the weight of the data. According to the measurement errors $\epsilon$ of each data point of CFs, the weight $W$ of $i^{\rm th}$ CF point is set to $1/\epsilon_{i}$ for the iterative process. $C^{(t)}$ is the predicted CF at the $t^{\rm th}$ iteration:
$C_{i}^{(t)}=\sum_{j}K_{ij}S_{j}^{(t)}.$ (9)
Additionally, it is crucial to consider a moderate number of iterations and an appropriate number of bins to ensure the algorithm’s robustness.
The transformation matrix $K_{ij}$ is calculated by
@@
III Method test
The imaging of the source and the extraction of interaction parameters ($f_{0}$,$d_{0}$) can now be achieved simultaneously using this method. To test the algorithm, we start with a CF calculated from the given source distribution and known $pp$ interactions. The known source function is modeled as a 1D Gaussian distribution with a mean of 00 and a standard deviation of $\sigma=3.5~{}\rm fm$ in the radial direction, as illustrated by blue solid curve in Fig. 2 (b). For imaging purposes, the source function is uniformly divided into $M=20$ bins with a width of $1\rm fm$ over the range of 00 to $20~{}\rm fm$ in the radial direction, while the CF is divided into $N=50$ bins with a width of $4~{}\rm MeV$ over the range of 00 to $200~{}\rm MeV$. The transformation matrix is calculated using the LL model, incorporating Coulomb interaction, with the interaction parameters for $pp$ interactions specified as $f_{0}=7.8~{}\rm fm$ and $d_{0}=2.77~{}\rm fm$. The CF used for imaging is obtained by multiplying the transformation matrix with the source function, incorporating the corresponding error values for each point, as shown by the blue line with the purple error band in Fig. 2 (a). In order to mimic the real measurement, the fluctuation is introduced to each point according to the experimental uncertainty. By treating the value $C$ of each point in the CF as the mean of a Gaussian distribution and the associated error as its standard deviation, we can sample numerous CFs. One of these is randomly selected to serve as the central CF, represented by the symbols of hollow blue circle in Fig. 2 (a).
Each sampled CF is then iteratively solved using the imaging method. Since the exact values of ($f_{0}$,$d_{0}$), key parameters of the transformation matrix, are unknown, it is reasonable to generate different matrices by exploring all possible values of ($f_{0}$,$d_{0}$). In this study, we consider values of ($f_{0}$,$d_{0}$) within the range of 00 to $10~{}\rm fm$, with a step size of $0.1~{}\rm fm$. For each sampled CF, we solve it using these matrices and compute the corresponding $\chi^{2}$ values ($\text{NDF}=N$) which represents for the overall deviation. Here $\chi^{2}$ is defined as
$\chi^{2}=\sum_{i}^{N}\left(\frac{C_{i}^{P}-C_{i}}{\sigma_{i}}\right)^{2},$ (11)
where $C_{i}^{P}$ represents the value of the $i^{\rm th}$ point. The matrix that yields the minimum $\chi^{2}$ is considered the best fit to describe the CF, allowing us to simultaneously determine ($f_{0}$,$d_{0}$), and the imaged source function.
@@ III Method test
The initial source function for the iteration is all assumed to be a Gaussian distribution with $\sigma_{0}=3~{}\rm fm$. Then the inverse problem is iteratively solved from a initial trial of the SF, which is taken as a Gaussian distribution with $\mu=0$ center location and arbitrary deviation. The reconstructed SF ends close to the known one, as marked SF1 in  2 (b). In order to check the robustness of algorithm against the variation of the initial SF, we also test the second initial trial, where the center of the initial Gaussian is arbitrarily set at $\mu=3$ fm instead of zero peaking. With the iteration going on, the output SF gradually approaches to the known one, as illustrated by the dashed curves in Fig.  2 (b). The marked SF2 sits on top of SF1, evidencing the robustness of the iterative procedure. Fig. 2 (c) presents the closeness quantity $\chi^{2}$ over iteration times in the imaging process of the central CF. During the procedure, ($f_{0}$,$d_{0}$) parameters are optimized by minimum $\chi^{2}$ and the closeness between the output and the known CF is checked in every iteration till the convergence condition is reached, namely the change of $\chi^{2}$ between two successive iterations is less than $1\times 10^{-4}$ , which functions as a threshold to avoid noise amplification. Fig. 2 (d) shows the minimum $\chi^{2}$ distribution on the ($f_{0}$,$d_{0}$) plane. The optimized ($f_{0}$,$d_{0}$), marked by red cross, converge to the input one (yellow star). The consistency between the input values and the reconstructed ones in panel (a), (b) and (d) demonstrates the validity of the deblurring algorithm.
Once all sampled CFs have been solved, we can calculate the standard deviations of ($f_{0}$,$d_{0}$), and each bin of the imaged source function, which represent the uncertainties in the imaging results. The solution obtained from the selected central CF is considered the central value. The corresponding $\chi^{2}$ values for different matrices of the central CF are shown in Fig. 2 (d), where the red star marks the optimal ($f_{0}$,$d_{0}$). The final solution for ($f_{0}$,$d_{0}$) is
$f_{0}=(7.5\pm 0.2)\rm~{}fm,d_{0}=(2.9\pm 1.5)\rm~{}fm,$ (12)
and the imaged source function is represented by the hollow red circles in Fig. 2 (b). In Fig. 2 (a), the red solid line represent the CF reconstructed using imaging methods. In addition, due to the introduction of a $1~{}\rm fm$ cutoff when calculating the Coulomb-inclusive wave function using the LL model, in order to prevent numerical divergence during the solution process, the first point in the source function is unreliable.
IV Application to data
@@
Now we apply the method to the $pp$ and $\bar{p}\bar{p}$ CFs in Au+Au at $\sqrt{S_{\rm NN}}=200~{}\rm{GeV}$ [29], taken by STAR collaboration [7, 30] at the Relativistic Heavy Ion Collider (RHIC) [31, 32]. The experimental CFs are inclusive and contain the contributions from the residual CFs of $p\Lambda$ and $\Lambda\Lambda$. The residual contribution, which shall be subtracted, is derived by the difference between the two fitting curves with and without the feedback corrections shown in [29]. Specifically, the difference between the fitting results of inclusive CFs and $1+x_{pp(\bar{p}\bar{p})}[C_{pp(\bar{p}\bar{p})}-1]$ are subtracted based on the experimental data. The relative fraction of protons and antiprotons are estimated by the simulation of THMINATOR2 model under the global DCA cut condition (DCA$<2cm$). The values of $x_{pp}$ and $x_{\bar{p}\bar{p}}$ are 0.45 and 0.42, respectively. In addition, we suppose that there are no differences between inclusive CF and true $pp(\bar{p}\bar{p})$ CF when $q_{rel}>160\rm MeV$, which means we use the experimental data of $q_{rel}>160\rm MeV$ directly. Subsequently, by combining the values of $x_{pp(\bar{p}\bar{p})}$, the true $pp$ ($\bar{p}\bar{p}$) CFs can be obtained. The pure $pp$ and $\bar{p}\bar{p}$ CFs are presented in Fig. 4 (a) by the symbols, where the uncertainties are taken same as in the experiment for each data point.
In the calculation of the Kernel $K_{ij}$, one needs the momentum distribution of $p$ and $\bar{p}$. The transverse momentum $p_{\rm t}$ distributions are taken from the midrapidity range in the same colliding system [33] measured by PHENIX experiment [34, 35]. The longitudinal rapidity $y_{\rm l}$ as well as the azimuth in transverse direction are uniformly distributed. The cuts $0.4<p_{\rm t}<2.5\rm{GeV}/c$ and $|y_{\rm l}|<0.7$ are taken as same as STAR data where the CFs are measured. And then, the three-dimension momentum distributions can be obtained. Specifically, we sample the rapidity uniformly within $-0.7<y_{1}<0.7$ ($y_{1}=\frac{1}{2}ln(\frac{1+\beta_{z}}{1-\beta_{z}})$), the $p_{t}$ from the distributions reported by the PHENIX experiment satisfied with the cuts of STAR, and the azimuthal angle $\varphi$ uniformly between 0 and $2\pi$. From these, we obtain the momentum of one particle randomly,
$p_{x}=p_{t}cos\varphi,p_{y}=p_{t}sin\varphi,p_{z}=\pm\sqrt{\frac{\beta_{z}^{2}(p_{t}^{2}+m_{p(\bar{p})}^{2})}{1-\beta_{z}^{2}}}.$ (13)
Based on the momentum sampling method described above, we present the three-dimensional momentum distributions for both protons and antiprotons under the given conditions. The momentum values required to generate the matrix are then sampled from these distributions.
@@
As well noticed, the LL model provides relatively weak constraints on $d_{0}$ when analyzing experimental data. So the widely accepted theoretical and experimental approximation $d_{0}=2.8\rm~{}fm$ is fixed here for both $pp$ and $\bar{p}\bar{p}$ pairs. The convergence criterion is set to $1\times 10^{-2}$. The corresponding optimal values of $f_{0}$ obtained are as follows:
$f_{0(pp)}=(7.7\pm 0.2)\text{fm},\quad f_{0(\bar{p}\bar{p})}=(7.9\pm 0.3)\text{fm},$ (14)
@@
and the corresponding distributions of $f_{0}$ are shown in the Fig. 3. The $f_{0}$ values for $pp$ and $\bar{p}\bar{p}$ pairs are identical and consistent with both the theoretical calculations and the experimental results from STAR collaboration [7].
Figure 3: (Color Online) The distribution of the optimal $f_{0}$ values obtained from the imaging of each sampled pseudo-CFs. The red and blue lines represent the optimal $f_{0}$ distributions for $pp$ pairs and $\bar{p}\bar{p}$ pairs, respectively.
The converged CFs of $pp$ and $\bar{p}\bar{p}$ reproduce fairly the experimental data, as demonstrated by the solid and dashed curves in Fig. 4 (a), respectively. The results of the restored source imaging are plotted in panel (b) by red dots for $p$ and blue circles for $\bar{p}$. The red and blue lines are represented for the extracted radii of $p$ and $\bar{p}$ source by STAR, respectively. The extracted radii refer to the single-particle distribution in the source rest frame, which differs from the $S(\mathbf{r})$ defined in this work. In our case, $S(\mathbf{r})$ is a function of the relative distance between two particles. To facilitate comparison, the Gaussian source radii extracted by the STAR collaboration have been multiplied by $\sqrt{2}$. In order to estimate the uncertainty of each imaging point, numerous pseudo-CFs for $pp$ and $\bar{p}\bar{p}$ are sampled, smearing the content in each individual $q$ bin by Gaussian fluctuation, for which the center (standard deviation) is taken from the experimental value (uncertainty). Then the standard deviation of the reconstructed SFs are calculated and plotted on each data point bin-by-bin as the error bar on the SF.
@@
Figure 4: (Color Online) (a) The reconstructed CFs of $pp$ (red solid curve) and $\bar{p}\bar{p}$ (blue dashed curve) pairs in comparison to the experimental CFs [7] (red dots for $pp$ and blue circles for $\bar{p}\bar{p}$ pairs). (b) Reconstructed SFs for $p$ (red dots) and $\bar{p}$ (blue circles) in Au+Au at center of mass energy 200 GeV. The symbols are the results of direct imaging without Gaussian assumption. The SFs with Gaussian shape obtained in [7] are shown by the curves for reference.

### v3

# Imaging Freeze-out Sources and Extracting Strong Interaction Parameters in Relativistic Heavy-Ion Collisions

@@ Abstract
By combining femtoscopic interferometry with an optical deblurring algorithm, we present a novel method to image the source in heavy-ion collisions (HICs) while simultaneously determining the interaction strength between particle pairs. The spatial distribution of the emission source has been reconstructed for protons ($p$) and antiprotons ($\bar{p}$) from the respective $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au collisions at $\sqrt{S_{\rm NN}}=200$ GeV. Within experimental uncertainties, protons and antiprotons share the same freeze-out distribution showing higher density in the center compared to the widely assumed Gaussian shape. The results evidence the matter-antimatter symmetry in coordinate space at the moment of freeze-out before the nucleons are fully randomized in the collisions.
Source Imaging, Relativistic Heavy-Ion Collisions, Strong Interaction, Anti-proton
@@
Introduction - It is well established that matter and antimatter were created in equal amounts during the first moments after the Big Bang. However, the observable universe consists overwhelmingly of matter, posing a fundamental question in physics. This puzzling phenomenon is related to the profound symmetry breaking in nature and remains a challenge in nuclear and particle physics. On modern accelerators, heavy ion collisions (HICs) create conditions akin to those in the early universe. So after the success of observing the production of antimatter [1, 2, 3], people can further investigate the spatial evolution and decoupling of the matter and antimatter on the femtoscopic scale.
@@
Hanbury Brown and Twiss (HBT) intensity interferometry, originally developed to measure the angular radius of Sirius [4], has been extensively adapted in nuclear physics [5, 6, 7]. It enables the extraction of spatiotemporal characteristics of the emission source in HICs by analyzing particle-pair correlation functions (CFs) at small relative momenta [8, 9]. Additionally, the spatial information embedded in the CFs has been applied to investigate the exotic structure and rare decay mode of unstable nuclei [10, 11, 12, 13, 14]. In a basic picture, the CF is the result of the low-energy scattering between the particle pair emitted from the source. It carries the information of the interaction strength, defined by the scattering length $f_{0}$ and effective range $d_{0}$ between the pair, written as ($f_{0}$, $d_{0}$) in short.
@@
Thus, simultaneous inference of the source distribution and the interaction strength becomes feasible for either particle or antiparticle pairs. Based on the Lednický-Lyuboshitz (LL) model, STAR Collaboration has verified that protons and antiprotons exhibit nearly identical interaction properties, and the extracted source size parameters for $pp$ and $\bar{p}\bar{p}$ pairs were found to be similar [15]. Specifically, within experimental uncertainties, the parameters ($f_{0}$, $d_{0}$) are equivalent for $pp$ and $\bar{p}\bar{p}$ pairs, consistent with CPT symmetry, a fundamental principle in physical laws. The Gaussian source assumption has become widely adopted in correlation function analysis, from which significant results have been obtained in the studies of particle interactions and source distributions [16, 17, 18, 19].
@@
However, more studies start to notice the non Gaussian feature of the source. By Gaussian parameterization it is assumed that all the participants are totally randomized in space prior to the freeze-out instant, and the initial state as well as the evolution details of the colliding system are lost. This strong assumption is unfavorable in the context of various recent studies. For two examples, i) high energy HIC has been applied to infer the initial shape of the colliding nuclei [20, 21]. ii) The Ohm’s law is suggestively broken because of the strong field and extremely short evolution time [22]. In both cases it is of importance to probe the spatial evolution of the colliding system. In recent experiments, a similar deviation from Gaussian distributions has been observed in the freeze-out source distributions of particles, exhibiting long tails and Lévy-stable characteristics [23, 24, 25, 26, 27, 28]. This behavior suggests that hadronic evolution could be described by Lévy walks, where the movement patterns are governed by long-tailed random walks. This concept has already played a significant role in the natural sciences [29, 30] and is gradually gaining widespread attention in the field of heavy-ion collisions. Thus, it is crucial to obtain the true distribution without pre-assuming a specific form such as the Gaussian.
@@
Despite of the difficulty in solving inverse problem, the direct imaging of the source has been achieved without introducing the analytic parameterization [31, 32, 33]. More recently, the Bayessian-based deblurring method, known as Richardson-Lucy (RL) algorithm [34, 35] which was initially developed for optical imaging [36, 37], has gained prominence in nuclear physics [38, 39, 40, 41, 42]. Particularly, the RL algorithm has been applied in source imaging in HICs, highlighting the method’s effectiveness and robustness [39, 43]. In this letter, combining the RL deblurring algorithm and the LL model incorporating the pair interaction, we bring the realisation to solve the inverse problem of source imaging from the experimental CFs. The source distribution and ($f_{0}$, $d_{0}$) are determined simultaneously. The similarity of the source spatial distribution of $p$ and $\bar{p}$ emission has been demonstrated.
@@
Methodology - Defined as a function of the relative momentum $\mathbf{q}=\mathbf{p}_{1}-\mathbf{p}_{2}$ in the particle pair rest frame (PRF), the CF is given experimentally by $C(\mathbf{q})=\frac{P(\mathbf{p}_{1},\mathbf{p}_{2})}{P(\mathbf{p}_{1})P(\mathbf{p}_{2})}$, where $P(\mathbf{p}_{1})$ and $P(\mathbf{p}_{2})$ are the probabilities of detecting particles with momenta $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, respectively, and $P(\mathbf{p}_{1},\mathbf{p}_{2})$ is the joint probability of detecting both particles simultaneously. For uncorrelated emissions, $C=1$ is expected. In theoretical frameworks, the correlation function (CF) is computed using the Koonin-Pratt (KP) formula [44, 45]
@@
where $S$ is the source function (SF) related to the relative distance $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ of the particle-pair in the source rest frame, and $\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})$ is the wave function calculated from the relative momentum $\mathbf{q}$ and distance $\mathbf{r}^{\prime}$ in the PRF. Based on the LL model [46], the interaction strength between particle pair is characterized by ($f_{0}$, $d_{0}$) describing the strong interaction scattering amplitude in S-wave expansion under the effective range approximation.
@@
In optical deblurring, the measured distribution ($f(\nu)$) is related to the true photon distribution ($\mathcal{F}(\mu)$) through a conditional probability function ($P(\nu|\mu)$),
@@
From the mathematical similarity between the KP formula (1) and the optical deblurring equation (2), it is naturally suggested that the RL algorithm can be applied iteratively to solve the SF. By evaluating the converged $\chi^{2}$ value, one can identify the optimal ($f_{0}$, $d_{0}$) for the wave function that best describes the CF, thereby achieving both SF imaging and extraction of strong interaction parameters simultaneously. For the convenience of numerical implementation, the KP formula is discretized by dividing $C$ into $N$ bins and $S$ into $M$ bins ($N\geq M$). After discretization, the RL algorithm is used to iteratively solve for the distribution of the source function [39, 42]. The CF $C$ is expressed by $S$ through an $N\times M$ transformation matrix $K$ as $C_{i}=\sum_{j=1}^{M}K_{ij}S_{j}$. This transformation matrix is calculated by
@@
Model test -The imaging of the source and the extraction of interaction parameters ($f_{0}$,$d_{0}$) can now be achieved simultaneously using this method. To test the algorithm, we start with calculating the CF from the given source distribution and known $pp$ interactions with $f_{0}=7.8$ fm and $d_{0}=2.77$ fm. In order to mimic the real measurement, the fluctuation is introduced to each point according to the experimental uncertainty. Fig. 2 (a) shows the theoretic CF (thick solid curve) and the fluctuated CF (open circle). Then the inverse problem is iteratively solved from a initial trial of the SF, which is taken as a Gaussian distribution with $\mu=0$ center location and arbitrary deviation. The reconstructed SF ends close to the known one, as marked SF1 in Fig. 2 (b). In order to check the robustness of algorithm against the variation of the initial SF, we also test the second initial trial, where the center of the initial Gaussian is arbitrarily set at $\mu=3$ fm instead of zero peaking. With the iteration going on, the output SF gradually approaches to the known one, as illustrated by the dashed curves in Fig.  2 (b). The marked SF2 sits on top of SF1, evidencing the robustness of the iterative procedure. Fig. 2 (c) presents the closeness quantity $\chi^{2}$ over iteration times in the imaging process of the central CF. Here $\chi^{2}$ is defined as $\chi^{2}=\sum_{i}^{N}\left(\frac{C_{i}^{P}-C_{i}}{\sigma_{i}}\right)^{2}$, where $C_{i}^{P}$ represents the value of the $i^{\rm th}$ point. During the procedure, ($f_{0}$,$d_{0}$) parameters are optimized by minimum $\chi^{2}$ and the closeness between the output and the known CF is checked in every iteration till the convergence condition is reached, namely the change of $\chi^{2}$ between two successive iterations is less than $1\times 10^{-4}$ , which functions as a threshold to avoid noise amplification. Fig. 2 (d) shows the minimum $\chi^{2}$ distribution on the ($f_{0}$,$d_{0}$) plane. The optimized ($f_{0}$,$d_{0}$), marked by red cross, converge to the input one (yellow star). The consistency between the input values and the reconstructed ones in panel (a), (b) and (d) demonstrates the validity of the deblurring algorithm.
@@
Application to data - Now we apply the method to the $pp$ and $\bar{p}\bar{p}$ CFs in Au+Au at $\sqrt{S_{\rm NN}}=200~{}\rm{GeV}$ [47], taken by STAR collaboration [15, 48] at the Relativistic Heavy Ion Collider (RHIC) [49, 50]. The experimental CFs are inclusive and contain the contributions from the residual CFs of $p\Lambda$ and $\Lambda\Lambda$. The residual contribution, which shall be subtracted, is derived by the difference between the two fitting curves with and without the feedback corrections shown in [47]. The pure $pp$ and $\bar{p}\bar{p}$ CFs are presented in Fig. 3 (a) by the symbols, where the uncertainties are taken same as in the experiment for each data point.
@@
In the calculation of the Kernel $K_{ij}$, one needs the momentum distribution of $p$ and $\bar{p}$. The transverse momentum $p_{\rm t}$ distributions are taken from the midrapidity range in the same colliding system [51] measured by PHENIX experiment [52, 53]. The longitudinal rapidity $y_{\rm l}$ as well as the azimuth in transverse direction are uniformly distributed. The cuts $0.4<p_{\rm t}<2.5\rm{GeV}/c$ and $|y_{\rm l}|<0.7$ are taken as same as STAR data where the CFs are measured. As well noticed, the LL model provides relatively weak constraints on $d_{0}$ when analyzing experimental data. And the variation of $d_{0}$ brings correspondingly uncertainty to the source function to be extracted. So the widely accepted theoretical and experimental approximation $d_{0}=2.8\rm~{}fm$ is fixed here for both $pp$ and $\bar{p}\bar{p}$ pairs. The convergence criterion is set to $1\times 10^{-2}$. The corresponding optimal values of $f_{0}$ obtained are as follows:
$f_{0(pp)}=(7.7\pm 0.2)\text{fm},\quad f_{0(\bar{p}\bar{p})}=(7.9\pm 0.3)\text{fm}.$ (4)
@@
The $f_{0}$ values for $pp$ and $\bar{p}\bar{p}$ pairs are identical and consistent with both the theoretical calculations and the experimental results from STAR collaboration [15].
@@
Figure 3: (Color Online) (a) Reconstructed correlation functions (CFs) for $pp$ (red solid curve) and $\bar{p}\bar{p}$ (blue dashed curve) pairs, compared with experimental data from [15] (red dots for pp and blue circles for $\bar{p}\bar{p}$ pairs). (b) Reconstructed SFs for $p$ (red dots) and $\bar{p}$ (blue circles) in Au+Au at center of mass energy 200 GeV. The symbols are the results of direct imaging without Gaussian assumption. The SFs with Gaussian shape obtained in [15] are shown by the curves for reference.
The converged CFs of $pp$ and $\bar{p}\bar{p}$ reproduce fairly the experimental data, as demonstrated by the solid and dashed curves in Fig. 3 (a), respectively. The results of the restored source imaging are plotted in panel (b) by red dots for $p$ and blue circles for $\bar{p}$. In order to estimate the uncertainty of each imaging point, numerous pseudo-CFs for $pp$ and $\bar{p}\bar{p}$ are sampled, smearing the content in each individual $q$ bin by Gaussian fluctuation, for which the center (standard deviation) is taken from the experimental value (uncertainty). Then the standard deviation of the reconstructed SFs are calculated and plotted on each data point bin-by-bin as the error bar on the SF.